5. A two-tailed hypothesis test for a repeated-measures research design Aa Aa A graduate student is interested in whether journaling can affect grief and healing. For her study, she selects a random sample of 64 adults who have been widowed within the past 10 to 14 months. The subjects complete a battery of psychological questionnaires before and after spending two months journaling about their emotions Before the two months of journaling, the mean score on the impact of event scale avoidance subscale, which measures how much individuals consciously avoid thoughts and feelings associated with the loss of their spouses, was 15.4. After the two months of journaling, the mean score was 15.0. The mean of the differences between each persons pre- and post- scores was 0.4, with a standard deviation of the differences equal to 1.9. The graduate student has no presupposed assumptions about whether journaling can affect grief and healing, so she formulates the null and alternative hypotheses as Ho: HD 30 H1: HD 0 She uses a repeated-measures t test. Because the sample size is large, if the null hypothesis is true as an equality, the test statistic follows a t-distribution with n 1 64 1 63 degrees of freedom. t Distribution Degrees of Freedom 55 4.0 3.0 2.0 1.0 0.0 1.0 2.0 3.0 4.0A two-talled hypothesis test for a repeated-measures research design A graduate student is interested in whether journaling can affect grief and healing. For his study, he selects a randor sample of 64 adults who have been widowed within the past 10 to 14 months. The subjects complete a battery or psychological questtonnalres before and after spending two months Journaling about their emotions. Before the two months of journaling, the mean score on the impact of event scale-avoidance subscale, which measures how much individuals consciously avoid thoughts and feelings associated with the loss of their spouses, was 15.4. After the two months of journaling, the mean score was 0.4 lower, with a standard deviation of 1.9. The graduate student has no presupposed assumptions about whether Journaling can affect grief and healing, so he formulates the null and alternative hypotheses as: H0:mu 0=0 H1:mu 0 0 He uses a repeated-measures t test.Because the sample size is large,if the null hypothesis is true as an equality,the test statistic follows a t-distribution with n-1 =64-1=63 degrees of freedom. This is a test. Use the distribution tool to find the critical score(s) for the level of significance alpha = 0.1. The critical score(s) (the value(s) for that separates(s) the tall(s) from the main body of the distribution, forming the
Answer #1

Result:

A two-talled hypothesis test for a repeated-measures research design A graduate student is interested in whether journaling can affect grief and healing. For his study, he selects a randor sample of 64 adults who have been widowed within the past 10 to 14 months. The subjects complete a battery or psychological questtonnalres before and after spending two months Journaling about their emotions. Before the two months of journaling, the mean score on the impact of event scale-avoidance subscale, which measures how much individuals consciously avoid thoughts and feelings associated with the loss of their spouses, was 15.4. After the two months of journaling, the mean score was 0.4 lower, with a standard deviation of 1.9. The graduate student has no presupposed assumptions about whether Journaling can affect grief and healing, so he formulates the null and alternative hypotheses as: H0:mu 0=0 H1:mu 0 0 He uses a repeated-measures t test.Because the sample size is large,if the null hypothesis is true as an equality,the test statistic follows a t-distribution with n-1 =64-1=63 degrees of freedom.

This is a two tailed test.

Use the distribution tool to find the critical score(s) for the level of significance alpha = 0.1. The critical score(s) (the value(s) for   (-2.656, 2.656)

Standard error = 1.9/sqrt(64) = 0.2375

Test statistic = 1.6842

t statistic not in the critical region.

Null hypothesis is not rejected.

The graduate student cannot conclude that journaling affects grief.

This means support group has more consistent treatment effect.

95% CI of mean difference would be wider for original study.

She will be more likely to reject the null hypothesis.